OpenStudy (anonymous):

Evaluate 32+(6-2)•4-6 over 3? 1) 20 2) 23 3) 50 4) 26

OpenStudy (anonymous):

What do you think?

OpenStudy (ali2x2):

32 + (7 - 2) x 4 - 6/3 = 32 + 5 x 4 - 6/3 Next you deal with the division and multiplication 32 + 5 x 4 - 6/3 = 32 + 20 - 2 Finally, you deal with the addition and subtraction in order 32 + 20 - 2 = 52 - 2 = 50

OpenStudy (anonymous):

thanks ali